Sodium phosphate is added to a solution that conta
Answer #1

14)

The two reaction that are happening here are

2Na3PO4 + 3CaCl2\rightarrow 6NaCl + Ca3(PO4)2

Na3PO4 + Al(NO3)3 \rightarrow AlPO4 + 3NaNO3

Aluminum phosphate
AlPO₄(s) ⇄ Al³⁺(aq) + PO₄³⁻(aq)

Ksp = [Al³⁺]∙[PO₄³⁻]
with Ksp = 9.84×10⁻²¹

Calcium phosphate
Ca₃(PO₄)₂(s) ⇄ 3 Ca²⁺(aq) + 2 PO₄³⁻(aq)

Ksp = [Ca²⁺]³∙[PO₄³⁻]²
with Ksp = 2.07×10⁻³³
source: http://www4.ncsu.edu/~franzen/public_html/CH201/data/Solubility_Product_Constants.pdf

When you solve the equilibrium equation of each salt for [PO₄³⁻] you get the level of phosphate concentration at which the salt starts to precipitate:
- aluminum phosphate
[PO₄³⁻] = Ksp / [Al³⁺] = 9.84×10⁻²¹ / 0.0094 = 1.04×10⁻¹⁸ M
- calcium phosphate
[PO₄³⁻] = √(Ksp / [Ca²⁺]²) = √(2.07×10⁻³³ / 0.025³) = 1.15×10⁻¹⁴ M

Because [PO₄³⁻] is smaller for Al³⁺, aluminum phosphate precipitates first.

when calcium phosphate starts to precipitate the aluminum ion concentration is
[Al³⁺] = Ksp / [PO₄³⁻] = 9.84×10⁻²¹ / 1.15×10⁻¹⁴ = 8.55×10⁻⁷ M